# LLM4DSR: Leveraing Large Language Model for Denoising Sequential Recommendation

### v1

@@ 3.3. Replace Noise with Suggested Items
Instead of simply removing noise, it is more logical to replace it with an item that closely aligns with the user’s preferences. Replacing rather than simply deleting can mitigates the issue of information loss resulting from erroneous deletions. By leveraging the powerful reasoning capabilities and extensive knowledge of LLMs, we can enable the LLM to select an item that better matches the user’s historical interaction patterns to replace the noise sample.
Formally, for a sequence , if is identified as noise (i.e., ), we use the prompt to generate a suggested item, referred to as the oracle. Due to the hallucination phenomenon inherent in LLMs, the oracle may be a fake item that does not necessarily exist in the item set. Inspired by BIGRec (Bao et al., 2023a), we employ grounding techniques to align it with the item set by searching for the item whose embedding is most similar to that of the oracle.
where denotes the embedding of the item and denotes the embedding of the suggested item’s title generated by the LLM. The grounding operation maps the generated item oracle to the actual item with the most similar embedding.
@@ 4. Experiments
RQ3: How does LLM4DSR identify noise and complete the sequence?
@@ 4.1. Experimental Settings
Three real-world datasets with 10-core setting: Amazon Video Games, Amazon Toy and Games, and Amazon Movies 222https://jmcauley.ucsd.edu/data/amazon/index_2014.html were utilized in our experiments, which have been used in the previous works (Cui et al., 2024). To ensure a fair comparison, we adopted the data preprocessing used in recent studies (Bao et al., 2023a). For each user sequence longer than 11 interactions, a sliding window of length 11 is used to segment the sequences. The sequences were then organized in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1. We randomly retained 20,000 items for Amazon Movies due to its large size. The dataset statistics are presented in Table 1.
Two noise settings are used in the experiment. The first setting uses the original real-world datasets, which are assumed to inherently contain natural noise. The second setting involves artificial noise, where we randomly replace items in the user interaction sequences according to a specified probability. The noise ratio indicates that each item is randomly replaced with a probability of . Note that the artificial noise is introduced before segmenting the sequences using the sliding window, thus preventing any information leakage between sequences.
@@
We conducted experiments on three backbones, including the traditional sequential recommendation models SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019), as well as a LLM-based recommendation model LLaRA (Liao et al., 2024). The methods compared in the study fall into two categories:
Explicit Denoising Methods (HSD (Zhang et al., 2022), STEAM (Lin et al., 2023)): HSD identifies noise based on inconsistency signals, whereas STEAM employs traditional models to train a discriminator and generator for noise identification and completion. 333We acknowledge the existence of other explicit denoising methods, such as SSDRec (Zhang et al., 2024a) and END4Rec (Han et al., 2024). We excluded them from our study because it necessitates additional information, which is incompatible with our experimental setup. Note that we performed end-to-end training on the original backbone (i.e., BERT4Rec) for HSD and STEAM and after completing this training, we transferred the generated denoised dataset to other backbones for training from scratch.
Implicit Denoising Methods (FMLP (Zhou et al., 2022), CL4SRec (Xie et al., 2022)): FMLP introduces FFT to remove noise signals from the representation, and CL4SRec enhances sequences during contrastive learning to increase robustness against noise. We only test implicit denoising methods on the traditional sequential recommendation backbone due to the coupling of implicit denoising methods with specific models, which prevents their application to LLM-based recommendation systems.
@@
For training LLM4DSR, we select Llama3-8B (Dubey et al., 2024) as the backbone and randomly sample 5000 prompts to fine-tune the model over 50 epochs using LoRA (Hu et al., 2021). For the threshold hyperparameter , we conduct a search within the [0,1] quantiles of the noise probabilities across all items, with a search interval of 0.05. For traditional models, we fix the learning rate at 0.001 and the embedding size at 64. We use the source code provided in the original papers and search for optimal hyperparameters for all comparison methods according to the instructions in the original papers. The evaluation metrics employed were NDCG@ and Hit Rate@, with set to 20 and 50.
| Games | Toy | Movie | |||||||||||
| Backbone | Method | NDCG@20 | NDCG@50 | HR@20 | HR@50 | NDCG@20 | NDCG@50 | HR@20 | HR@50 | NDCG@20 | NDCG@50 | HR@20 | HR@50 |
| None | 0.0299 | 0.0377 | 0.0646 | 0.1040 | 0.0190 | 0.0239 | 0.0501 | 0.0749 | 0.0468 | 0.0548 | 0.0900 | 0.1308 | |
@@
| HSD | 0.0246 | 0.0324 | 0.0514 | 0.0908 | 0.0121 | 0.0172 | 0.0319 | 0.0582 | 0.0369 | 0.0451 | 0.0780 | 0.1196 | |
@@
| FMLP | 0.0369 | 0.0449 | 0.0710 | 0.1116 | 0.0195 | 0.0242 | 0.0426 | 0.0699 | 0.0554 | 0.0641 | 0.0952 | 0.1394 | |
@@
[…]
| LLaRA | LLM4DSR | 0.0217 | 0.0263 | 0.0352 | 0.0588 | 0.0131 | 0.0169 | 0.0265 | 0.0455 | 0.0166 | 0.0193 | 0.0258 | 0.0392 |
| Backbone | Method | Games | Toy | Movie | |||||||||
| NDCG@20 | NDCG@50 | HR@20 | HR@50 | NDCG@20 | NDCG@50 | HR@20 | HR@50 | NDCG@20 | NDCG@50 | HR@20 | HR@50 | ||
| SASRec | None | 0.0252 | 0.0314 | 0.0526 | 0.0838 | 0.0139 | 0.0178 | 0.0348 | 0.0542 | 0.0390 | 0.0455 | 0.0750 | 0.1084 |
[…]
| HSD | 0.0198 | 0.0250 | 0.0412 | 0.0678 | 0.0090 | 0.0127 | 0.0254 | 0.0444 | 0.0340 | 0.0407 | 0.0700 | 0.1042 | |
[…]
| LLM4DSR | 0.0292 | 0.0357 | 0.0624 | 0.0956 | 0.0152 | 0.0200 | 0.0401 | 0.0642 | 0.0431 | 0.0517 | 0.0876 | 0.1314 | |
| BERT4Rec | None | 0.0266 | 0.0335 | 0.0534 | 0.0880 | 0.0150 | 0.0195 | 0.0346 | 0.0574 | 0.0485 | 0.0569 | 0.0826 | 0.1258 |
| FMLP | 0.0281 | 0.0353 | 0.0598 | 0.0962 | 0.0147 | 0.0188 | 0.0330 | 0.0536 | 0.0478 | 0.0561 | 0.0818 | 0.1240 | |
[…]
| LLM4DSR | 0.0316 | 0.0397 | 0.0644 | 0.1056 | 0.0183 | 0.0231 | 0.0401 | 0.0645 | 0.0509 | 0.0591 | 0.0890 | 0.1306 | |
[…]
| STEAM | 0.0108 | 0.0146 | 0.0224 | 0.0416 | 0.0092 | 0.0120 | 0.0202 | 0.0348 | 0.0107 | 0.0127 | 0.0182 | 0.0286 | |
| LLM4DSR | 0.0150 | 0.0193 | 0.0286 | 0.0504 | 0.0121 | 0.0162 | 0.0280 | 0.0482 | 0.0159 | 0.0186 | 0.0250 | 0.0390 | |
@@ 4.2. Performance Comparison (RQ1)
In this section, we analyze the superior of LLM4DSR under two different noise setting as compared with other baselines.
4.2.1. Evaluations on Real-world Datasets.
In Table 2, we present the test results of various methods on real-world datasets. Notably, LLM4DSR consistently achieved the best performance across almost all datasets.
For explicit denoising methods (i.e., STEAM and HSD), we observed that while these methods perform well on BERT4Rec, their effectiveness diminishes when retraining on SASRec using the denoised dataset they generated. For instance, on the Games dataset, the STEAM method improves NDCG@20 from 0.0351 to 0.0373 on BERT4Rec. However, when the same denoised dataset is evaluated on SASRec, NDCG@20 drops from 0.0299 to 0.0233. This suggests that the denoised dataset generated by STEAM lacks good transferability. Additionally, the performance of HSD significantly declined. We argue that this issue may arise from HSD’s inability to effectively capture sequence information in short sequence scenarios. This limitation can result in the loss of critical information when incorrect items are deleted. LLM4DSR also demonstrated impressive performance when applied to LLM-based architectures (i.e., LLaRA). Notably, HSD demonstrates better performance on LLM-based backbones compared to traditional models. This may be due to the fact that the shortened sequences generated by HSD are more compatible with LLMs (Hou et al., 2024), and the inherent reasoning capabilities of LLMs make them relatively insensitive to erroneous deletions.
@@
4.2.2. Evaluations on Artificially Noisy Dataset.
The test results for various methods on the artificially noise-added dataset, with a noise ratio of 10%, are presented in Table 3. These results indicate that LLM4DSR consistently outperforms all comparative methods. Artificial noise, being more distinguishable than natural noise in real-world datasets, makes the improvement of LLM4DSR more pronounced on such datasets. For example, on the Game dataset, LLM4DSR combined with SASRec achieves a 5.3% improvement in NDCG@20, whereas on the artificially noise-added dataset, the improvement reaches 15.8%. Furthermore, we evaluated the performance of all comparative methods under different noise ratios, as illustrated in Figure 4. LLM4DSR significantly surpasses other comparative methods demonstrate its robust performance across varying noise conditions.
@@ 4.3. Ablation Study (RQ2)
To verify the contributions of various components of LLM4DSR, we conducted an ablation study on different variants under SASRec, using both real and artificially noisy datasets from the Game and Toy datasets. The variants tested include: (1) ”w/o uncertainty,” which omits uncertainty estimation and directly uses the LLM output for denoising; (2) ”w/o completion,” which excludes the completion capability of LLM4DSR and performs only deletion operations on sequences; and (3) ”w/ popularity completion,” which replaces noise with items randomly sampled from the item set based on popularity weights. As presented in Table 4, the consistently superior performance of LLM4DSR across all datasets highlights the effectiveness of each component. In particular, the performance decline observed in the ”w/o completion” and ”w/ popularity completion” emphasize the critical role of the replacement operation, demonstrating that the LLM4DSR can effectively extract user interests and provide items that better align with the sequence.
@@
4.4. Case Study (RQ3)
In this case study, we illustrate how LLM4DSR leverages open knowledge to achieve denoising. As depicted in Figure 6, we analyze a user’s viewing history from the Movie dataset, which predominantly consists of films in the drama, romance, and comedy genres. A notable outlier in this collection is the science fiction film ”Escape from the Planet of the Apes”. Utilizing prior knowledge about movie classifications, the LLM identifies this film as noise with a probability of . To enhance the sequence information, LLM4DSR recommends ”The Man from Snowy River”, a film that aligns with the user’s preferred drama/romance genres, as a suitable replacement for the identified noise.

### v2

@@ 3.3. Replace Noise with Suggested Items
Beyond removing noise items, we propose leveraging the generative capabilities of LLMs to suggest replacement items. The replacement items could bring additional knowledge of user preferences, potentially leading to improved recommendation performance. Specifically, for each identified noisy item , we use the prompt <User Interactions, Noise Items: , Suggested Items: > to guide the LLM in generating a description of the suggested replacement item . Given the hallucination issue of LLMs, could be a fake item that does not necessarily exist in the item set. Grounding techniques (Bao et al., 2023a) can be applied to find the most close existing item:
where represents the similarity between two textual descriptions, which can be implemented via the embedding similarity of texts. The item is then used to replace , generating a corrected sequence that can be leveraged to enhance the performance of various recommendation models.
@@ 4. Experiments
RQ3: How does hyperparameter affect model performance?
RQ4: How does LLM4DSR identify noise and complete the sequence?
@@ 4.1. Experimental Settings
Three conventional real-world datasets: Amazon Video Games, Amazon Toy and Games, and Amazon Movies 333https://jmcauley.ucsd.edu/data/amazon/index_2014.html were utilized in our experiments. These datasets are commonly used for the studies of LLM-based sequential recommendation (Cui et al., 2024; Bao et al., 2023a; Cao et al., 2024; Lee et al., 2024). To ensure a fair comparison, we adopted the same data preprocessing used in recent studies (Bao et al., 2023a). That is, for each user sequence longer than 11 interactions, a sliding window of length 11 is used to segment the sequences. The sequences were then organized in ascending order of timestamps to partition each dataset into training, validation, and testing sets with ratios of 8:1:1. We randomly retained 20,000 items for Amazon Movies due to its large size. The dataset statistics are presented in Table 2.
We employ two noise settings in our experiments: (1) the natural noise setting, which directly uses the original real-world datasets, as these datasets inherently contain noise introduced during practical data collection; and (2) the artificial noise setting, where a certain ratio of noise, denoted by , is explicitly injected into the dataset. Specifically, in the artificial noise setting, each item in a sequence has a probability of of being replaced by a randomly selected noise item. This setting allows us to evaluate the effectiveness of the denoising strategy under varying levels of noise. Moreover, the artificial noise setting provides ground-truth labels for the noise items, which can be used to assess the accuracy of different methods in identifying noise items. Note that the artificial noise is introduced before segmenting the sequences using the sliding window, thus preventing any information leakage between sequences.
@@
We conducted experiments on three backbones, including the traditional sequential recommendation models SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019), as well as a LLM-based recommendation model LLaRA (Liao et al., 2024). We compare LLM4DSR with the following denoising baselines 444We acknowledge the existence of other explicit denoising methods for sequential recommendation, such as SSDRec (Zhang et al., 2024a), END4Rec (Han et al., 2024). We excluded them from our study because these methods are not compatible with the setting of this paper.:
HSD (Zhang et al., 2022) is an explicit denoising method. It posits that noisy interactions within a sequence do not align with the user’s long-term and short-term interests and exhibit noticeable discontinuities with the sequence context. It jointly identifies noise items in the sequence from both user-level and sequence-level signals.
STEAM (Lin et al., 2023) is an explicit denoising method. It constructs a self-supervised task by randomly replacing items in the original sequence, training a discriminator and a generator to determine whether an item in the sequence is noise and to complete the sequence, respectively.
FMLP (Zhou et al., 2022) is an implicit denoising method. It uses Fast Fourier Transform and a learnable filter to eliminate noise in the representation.
CL4SRec (Xie et al., 2022) is an implicit denoising method. It constructs positive samples of the original sequence through transformations such as item cropping, item masking, and item reordering and then uses contrastive learning to enhance the model’s robustness to noise in the sequence.
We tested implicit denoising methods only on traditional sequential recommendation backbones. This is due to the coupling of these methods with specific models, which prevents their application to LLM-based recommendation systems.
4.1.3. Evaluation Metrics.
We utilize the following metrics to evaluate our method. The evaluation metrics employed were NDCG@ and Hit Rate@, with set to 20 and 50.:
Hit Ratio@K measures the proportion of positive items that appear in the top K positions of the recommendation list.
(5) where denotes an indicator function; denotes the positive item set in the test data for a user ; denotes the ranking position of the item for the user .
NDCG@K measures the ranking quality of recommendation through discounted importance based on the position:
(6) is the value of the ideal ranking with the optimal ranking for the user .
@@
For training LLM4DSR, we select Llama3-8B (Dubey et al., 2024) as the backbone and randomly sample 5000 prompts to fine-tune the model over 50 epochs using LoRA (Hu et al., 2021). For the threshold hyperparameter , we conduct a search within the [0,1] quantiles of the noise probabilities across all items, with a search interval of 0.05. In the experiment, we found that demonstrated strong robustness. Searching above the 0.7 quantile yielded satisfying results. A sensitivity analysis of this parameter is provided in Section 4.4. Additionally, to prevent excessive modification of any sequence, we limit each sequence to be modified by at most samples where the parameter is chosen from .
For SASRec and BERT4Rec, we fix the learning rate at 0.001 and the embedding size at 64. For the LLM-based recommendation model, LLaRA, the original configuration involves selecting the Top 1 item from the candidate items as the recommendation result, without item ranking process, which prevents the calculation of previously introduced evaluation metrics. To align with standard experimental settings, we follow (Bao et al., 2023a) and implement a grounding operation to enable LLaRA rank the entire item set.
We use the source code provided in the original papers and search for optimal hyperparameters for all comparison methods according to the instructions in the original papers. Besides, for the two explicit denoising methods, HSD and STEAM, we performe end-to-end training on the backbone used in the original paper (i.e., BERT4Rec) and generate the denoised dataset. To facilitate comparison with other backbones (i.e., SASRec and LLaRA), we transferred the denoised dataset to these backbones and trained them from scratch.
| Backbone | Method | Games | Toy | Movie | |||
| NDCG@20 | NDCG@50 | NDCG@20 | NDCG@50 | NDCG@20 | NDCG@50 | ||
| SASRec | None | 0.0299 | 0.0377 | 0.0190 | 0.0239 | 0.0468 | 0.0548 |
| FMLP | 0.0297 | 0.0377 | 0.0200 | 0.0245 | 0.0488 | 0.0574 | |
[…]
| HSD | 0.0246 | 0.0324 | 0.0121 | 0.0172 | 0.0369 | 0.0451 | |
| STEAM | 0.0233 | 0.0325 | 0.0187 | 0.0242 | 0.0475 | 0.0569 | |
@@
| FMLP | 0.0369 | 0.0449 | 0.0195 | 0.0242 | 0.0554 | 0.0641 | |
[…]
| HSD | 0.0301 | 0.0387 | 0.0187 | 0.0230 | 0.0467 | 0.0565 | |
[…]
| LLM4SRD | 0.0375 | 0.0464 | 0.0222 | 0.0280 | 0.0604 | 0.0697 | |
[…]
| HSD | 0.0231 | 0.0305 | 0.0136 | 0.0188 | 0.0271 | 0.0347 | |
[…]
| LLM4SRD | 0.0290 | 0.0375 | 0.0241 | 0.0306 | 0.0409 | 0.0504 | |
| Backbone | Method | Games | Toy | Movie | |||
| HR@20 | HR@50 | HR@20 | HR@50 | HR@20 | HR@50 | ||
| SASRec | None | 0.0646 | 0.1040 | 0.0501 | 0.0749 | 0.0900 | 0.1308 |
[…]
| HSD | 0.0514 | 0.0908 | 0.0319 | 0.0582 | 0.0780 | 0.1196 | |
[…]
| LLM4SRD | 0.0696 | 0.1148 | 0.0557 | 0.0830 | 0.0976 | 0.1440 | |
[…]
| FMLP | 0.0710 | 0.1116 | 0.0426 | 0.0699 | 0.0952 | 0.1394 | |
@@
| LLM4SRD | 0.0748 | 0.1202 | 0.0505 | 0.0799 | 0.1046 | 0.1516 | |
[…]
| HSD | 0.0490 | 0.0866 | 0.0357 | 0.0617 | 0.0688 | 0.1072 | |
[…]
| LLM4SRD | 0.0602 | 0.1026 | 0.0576 | 0.0905 | 0.0870 | 0.1350 | |
| Backbone | Method | Games | Toy | Movie | |||
| NDCG@20 | NDCG@50 | NDCG@20 | NDCG@50 | NDCG@20 | NDCG@50 | ||
| SASRec | None | 0.0252 | 0.0314 | 0.0139 | 0.0178 | 0.0390 | 0.0455 |
[…]
| HSD | 0.0198 | 0.0250 | 0.0090 | 0.0127 | 0.0340 | 0.0407 | |
| STEAM | 0.0261 | 0.0333 | 0.0123 | 0.0161 | 0.0401 | 0.0476 | |
| LLM4SRD | 0.0292 | 0.0357 | 0.0152 | 0.0200 | 0.0431 | 0.0517 | |
| BERT4Rec | None | 0.0266 | 0.0335 | 0.0150 | 0.0195 | 0.0485 | 0.0569 |
| FMLP | 0.0281 | 0.0353 | 0.0147 | 0.0188 | 0.0478 | 0.0561 | |
[…]
| HSD | 0.0288 | 0.0361 | 0.0152 | 0.0189 | 0.0390 | 0.0459 | |
[…]
| LLM4SRD | 0.0316 | 0.0397 | 0.0183 | 0.0231 | 0.0509 | 0.0591 | |
[…]
| HSD | 0.0146 | 0.0198 | 0.0116 | 0.0153 | 0.0197 | 0.0267 | |
[…]
| LLM4SRD | 0.0252 | 0.0321 | 0.0150 | 0.0204 | 0.0358 | 0.0445 | |
| Backbone | Method | Games | Toy | Movie | |||
| HR@20 | HR@50 | HR@20 | HR@50 | HR@20 | HR@50 | ||
| SASRec | None | 0.0526 | 0.0838 | 0.0348 | 0.0542 | 0.0750 | 0.1084 |
[…]
| HSD | 0.0412 | 0.0678 | 0.0254 | 0.0444 | 0.0700 | 0.1042 | |
[…]
| LLM4SRD | 0.0624 | 0.0956 | 0.0401 | 0.0642 | 0.0876 | 0.1314 | |
| BERT4Rec | None | 0.0534 | 0.0880 | 0.0346 | 0.0574 | 0.0826 | 0.1258 |
| FMLP | 0.0598 | 0.0962 | 0.0330 | 0.0536 | 0.0818 | 0.1240 | |
@@
| LLM4SRD | 0.0514 | 0.0866 | 0.0380 | 0.0655 | 0.0770 | 0.1216 | |
@@ 4.2. Performance Comparison (RQ1)
In this section, we analyze the superior of LLM4DSR under two different noise settings as compared with other baselines.
4.2.1. Evaluations on the nature noise setting.
In Table 3 and Table 4, we present the test results of various methods on real-world datasets. Notably, LLM4DSR consistently achieves the best performance across almost all datasets.
For explicit denoising methods (i.e., STEAM and HSD), we observed that while these methods perform well on BERT4Rec, their effectiveness diminishes when retraining on SASRec using the denoised dataset they generated. For instance, on Amazon Games, STEAM improves NDCG@20 from 0.0351 to 0.0373 on BERT4Rec. However, when the same denoised dataset is evaluated on SASRec, NDCG@20 drops from 0.0299 to 0.0233. This suggests that the denoised dataset generated by STEAM lacks good transferability. Additionally, the performance of HSD significantly declined. We argue that this issue may arise from HSD’s inability to effectively capture sequence information in short sequence scenarios. This limitation can result in the loss of critical information when incorrect items are deleted. In contrast, LLM4DSR demonstrates consistently superior performance across all datasets and backbones, proving its ability to effectively identify and replace noisy samples based on the internal knowledge of LLMs.
@@
4.2.2. Evaluations on the artifically noise setting.
The test results for various methods on the artifically noise setting, with a noise ratio of 10%, are presented in Table 5 and Table 6. The results indicate that LLM4DSR consistently outperforms all comparative methods. Artificial noise, being more distinguishable than natural noise in conventional real-world datasets, makes the improvement of LLM4DSR more pronounced on such datasets. For example, on Amazon Games, LLM4DSR combined with SASRec achieves a 5.3% improvement in NDCG@20, whereas on the artificially noise-added dataset, the improvement reaches 15.8%. Furthermore, we evaluated the performance of all comparative methods under different noise ratios, as illustrated in Figure 6. LLM4DSR significantly surpasses other comparative methods demonstrate its robust performance across varying noise conditions.
@@ 4.3. Ablation Study (RQ2)
4.3.1. The impact of different components in LLM4DSR
To verify the contributions of various components of LLM4DSR, we conducted an ablation study on different variants under SASRec, using both real and artificially noisy datasets from the Game and Toy datasets. The variants tested include: (1) ”w/o uncertainty,” which omits uncertainty estimation and directly uses the LLM output for denoising; (2) ”w/o completion,” which excludes the completion capability of LLM4DSR and performs only deletion operations on sequences; and (3) ”w/ popularity completion,” which replaces noise with items randomly sampled from the item set based on popularity weights. As presented in Table 7, the consistently superior performance of LLM4DSR across all datasets highlights the effectiveness of each component. In particular, the performance decline observed in the ”w/o completion” and ”w/ popularity completion” highlights the critical role of the replacement operation. This demonstrates that LLM4DSR effectively extracts user interests and suggests items that better align with the sequence, while also mitigating the negative impact of erroneous deletions.
4.3.2. Performance of separately denoising on training and test sets.
To more effectively demonstrate the impact of LLM4DSR on model performance, we evaluated the model after denoising the training and testing sets separately. As depicted in Figure 8, we conducted experiments on the Game and Toy datasets. Figure 8(a) presents the results on real-world datasets, while Figure 8(b) shows results on datasets with 10% artificial noise. Across all datasets and settings, denoising both the training and testing sets individually enhanced performance, with greater improvements observed when denoising the training set. Furthermore, simultaneously denoising both sets led to the most significant performance gains. The results highlight the importance of clean data in both training and evaluation phases, and demonstrate that LLM4DSR can be a valuable tool in enhancing data quality, ultimately leading to better model performance.
@@ 4.3.1. The impact of different components in LLM4DSR
4.4. Sensitivity analysis of hyperparameter (RQ3)
In the Uncertainty Estimation module, serves as the threshold distinguishing classification noise from clean samples. As increases, fewer items are identified as noise, leading to a lower modification ratio of the original sequence. To better understand the impact of on model performance, we examined the relationship between the hyperparameter and model performance under two noise settings, as illustrated in Figure 9. The x-axis represents the quantile of ; a quantile of 1 indicates that all samples are classified as clean, equivalent to training on the original dataset, whereas a quantile of 0 indicates that all samples are identified as noise and replaced (note that we limit the maximum number of modified items per sequence to ). We adjusted only for the training set, without filtering noise from the test data, to accurately reflect the impact of parameter changes. Figure 9(a) shows the results on the natural noise setting, while Figure 9(b) presents the results on the artificial noise setting.
In the artificially noisy Game dataset, NDCG@20 initially increases and then decreases as decreases. During the initial increase, LLM4DSR effectively filters out a large number of high-confidence noise samples; subsequently, performance declines due to modifications of some clean samples. In the real-world game dataset, NDCG@20 shows a second drop, which may be due to the fact that a large number of noise samples are concentrated in the samples with medium confidence. In the Movie, NDCG@20 increases with a decrease in and then remains stable, suggesting that LLM’s replacement operations effectively compensate for the information loss caused by erroneous deletions, thus enhancing the robustness of the parameter. Overall, is relatively insensitive to adjustments, and satisfactory results can be achieved by searching within a larger range (e.g., above the 0.7 quantile) during the actual application.
4.5. Case Study (RQ4)
In this case study, we illustrate how LLM4DSR leverages open knowledge to achieve denoising. As depicted in Figure 10, we analyze a user’s viewing history from the Movie dataset, which predominantly consists of films in the drama, romance, and comedy genres. A notable outlier in this collection is the science fiction film ”Escape from the Planet of the Apes”. Utilizing prior knowledge about movie classifications, the LLM identifies this film as noise with a probability of . To enhance the sequence information, LLM4DSR recommends ”The Man from Snowy River”, a film that aligns with the user’s preferred drama/romance genres, as a suitable replacement for the identified noise. Traditional ID-based denoising methods encounter substantial challenges in identifying noise due to their lack of access to semantic and categorical information about items, as well as the absence of supervisory signals from noisy labels. Moreover, these models do not possess reasoning and comprehension capabilities, further complicating the process of noise identification.